OpenStudy (anonymous):

COULD ANYONE HELP ME WITH MY HOMEWORK I HAVE A BIG TEST TOMORROW IN GEOMETRY CLASS

OpenStudy (anonymous):

hi there

OpenStudy (anonymous):

HELLO

OpenStudy (anonymous):

how may i help you..

OpenStudy (anonymous):

I NEED HELP WITH MY HOMEWORK IT GEOMETRY

OpenStudy (anonymous):

What is your equation?

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):



OpenStudy (anonymous):

Use trig, the 90 degree angle will tell you the length of the opposite side, and the use pythagorean theorem a^2 + b^2 = c^2 currently you have 10^2 + b^2 = c^2 b^2 = x c^2 = hypotenuse.

OpenStudy (anonymous):

yepp